OpenStudy (anonymous):

Write three fractions that are equivalent to 6/9 Show your work

OpenStudy (zaarin):

6*3/9*3=18/27 6*2/9*2=12/18 6*9/9*9=54/81

OpenStudy (anonymous):

For this u can * this in all the numbers u like ! Just this :)

OpenStudy (anonymous):

Thanxs for help

OpenStudy (anonymous):

Do u have any problem?:)

OpenStudy (anonymous):

2/3 4/6 16/24